# Cointegration with Occasionally Binding Constraints

### v3

@@ 3.3.2 Case (ii): kinked cointegration
{adjustwidth}
-2cm-2cm Refer to caption Refer to caption Refer to caption Refer to caption
@@
This case is perhaps more reminiscent of linear cointegration, which it accommodates as a special case. Here $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$ trace out ‘half’ subspaces of the same dimension $q$, but they need not be parallel, giving rise to a kink in $\mathscr{M}$ at the origin, with $\mathscr{M}$ itself being a linear cone. In general, $(Y,X)$ will follow a kinked Brownian motion driven by $p$ linear BMs, whose loadings, and the associated $r$ cointegrating relations, vary with the sign of $Y$. In the top panel of 3.3, $y_{t}$ and $x_{t}$ are cointegrated, but with distinct cointegrating vectors $\beta^{+}=(1,-1)^{\mathsf{T}}$ or $\beta^{-}=(\tfrac{1}{2},-1)^{\mathsf{T}}$ applying on $\mathscr{Z}^{+}$ or $\mathscr{Z}^{-}$, i.e. when $y_{t}$ is positive or negative. We have a kind of ‘threshold cointegration’, with the movement of $y_{t}$ across zero causing the model to switch between distinct cointegrating spaces (cf. Saikkonen and Choi, 2004). This switch is reflected in the trajectories of $y_{t}$ and $x_{t}$: when $y_{t}\geq 0$, the two series move together approximately one-for-one; whereas if $y_{t}\leq 0$, $y_{t}$ changes by two units for every one-unit change in $x_{t}$, in the long run. Relative to case (i), the trajectories of $\{y_{t}\}$ will now much more closely resemble those of a linear unit root process, and $\{y_{t}\}$ will accordingly tend to spend long stretches in both the positive and negative regions, with no tendency to revert to one or the other.
The bottom panel of 3.3 presents an important special case where the cointegrating relationships are such that $\{x_{t}\}$ behaves like an $I^{\ast}(0)$ process when $y_{t}<0$, but cointegrates with $y_{t}$ when $y_{t}>0$; the limiting process $X$ will thus be a censored Brownian motion.
@@
Suppose that $\psi=1$, so that this model (with $\overline{\pi}_{t}=\overline{\pi}=0$) may be rendered as
@@
$r_{t}^{\ast}$ $=r_{t-1}^{\ast}+\eta_{t},$ (2.6a)
and (LABEL:natrate21a) implies that the natural rate $r_{t}^{\ast}$ follows a random walk. Then similarly to (LABEL:PIpmexample), we have from the corresponding CKSVAR for $(i_{t},\pi_{t})$, as given in (LABEL:cksvar-natrate) above, that
@@
where $\gamma^{-1}\tau_{\mu}=\theta(1-\mu)(1-\theta\gamma)^{-1}\leq 0$, with strict inequality unless $\mu=1$. Thus $\operatorname{rk}\Pi^{+}=\operatorname{rk}\Pi^{-}=1$, and (by 4.2 below), there is a common stochastic trend present in both regimes – but when $i_{t}>0$ this loads only on $i_{t}$, and the cointegrating vector is $\beta^{+}=(0,1)^{\mathsf{T}}$. If unconventional policy is as effective as conventional policy ($\mu=1$), this holds also when $i_{t}<0$; otherwise the trend is shared by both $i_{t}$ and $\pi_{t}$, and their cointegrating vector in the negative regime is $\beta^{-}=(\gamma^{-1}\tau_{\mu},1)^{\mathsf{T}}$. Qualitatively, the behaviour of the series is as plotted for $(y_{t},x_{t})$ in the second row of 3.3 (with $y_{t}=-i_{t}$ and $x_{t}=-\pi_{t}$, i.e. with the signs reversed); in the limit, $n^{-1/2}\pi_{\lfloor n\lambda\rfloor}$ will be a censored BM (from above).
To account for this in economic terms, recall that (LABEL:is-pc21a) and (LABEL:taylor21a) imply that when $i_{t}>0$, the central bank is able to fully stabilise inflation, in the sense that $\pi_{t}=\overline{\pi}_{t}+(1-\gamma\theta)^{-1}\varepsilon_{t}=(1-\gamma\theta)^{-1}\varepsilon_{t}$, since $\overline{\pi}_{t}$ is constant and normalised to zero. Thus $\pi_{t}\mathbf{1}\{i_{t}>0\}\sim I^{\ast}(0)$; whereas, $i_{t}^{+}\sim I^{\ast}(1)$, since it inherits the stochastic trend in the natural rate of interest. However, if the model solution requires $i_{t}<0$, then the ZLB constraint inhibits the operation of monetary policy (if $\mu<1$), and inflation begins to drift away from its target. That drift corresponds to the stochastic trend in $r_{t}^{\ast}$, which is thus present in both $i_{t}^{-}$ and $\pi_{t}$, and hence these series cointegrate.
By contrast, if monetary policy is not effectively constrained by the ZLB, then $\pi_{t}$ would remain $I^{\ast}(0)$, irrespective of the sign of $i_{t}$. Thus the long run behaviour of $\pi_{t}$ – whether it is $I^{\ast}(0)$, or whether it follows a stochastic trend (and so is $I^{\ast}(1)$) when interest rates are at the ZLB – here provides identifying information on the relative effectiveness of unconventional monetary policy, i.e. on whether the ZLB is ever a truly binding constraint on the central bank, just as it did in 2.1.
@@
Suppose that $\delta=0$ in (LABEL:infltarget), so the model is now described by (LABEL:is-pc21b), (LABEL:taylor21b) and
@@
The inflation target thus remains time-varying, but there is no longer any feedback from past failures to hit that target. Then (LABEL:PIpmexample) simplifies to
@@
where $\varphi_{1}^{-1}\varphi_{\mu}=1+\frac{\theta(\gamma-1)(1-\mu)}{1-\theta\gamma}\in(0,1]$. Thus $i_{t}$ and $\pi_{t}$ are $I^{\ast}(1)$ and cointegrated, but with cointegrating relations $\beta^{+}=(1,-1)^{\mathsf{T}}$ and $\beta^{-}=(\varphi_{1}^{-1}\varphi_{\mu},-1)^{\mathsf{T}}$ that depend on the sign of $i_{t}$, unless $\mu=1$. Even though $i_{t}^{-}$ is unobserved, we can still distinguish between the cases $\mu=1$ and $\mu<1$ on the basis that, when $\mu<1$, the long-run variance of $\pi_{t}$ will differ depending on whether $i_{t}>0$ or $i_{t}=0$, as is evident in the behaviour of $x_{t}$ in the top panel of 3.3.8884.2 below implies that $n^{-1/2}\pi_{\lfloor n\lambda\rfloor}$ converges weakly to $\sigma_{\varepsilon}(\gamma-1)[1+(\varphi_{1}^{-1}\varphi_{\mu}-1)\mathbf{1}\{W(\lambda)<0\}]W(\lambda)$, for $\sigma_{\varepsilon}^{2}\coloneqq\mathbb{E}\varepsilon_{t}^{2}$ and $W$ a standard BM.
@@ 3.3.3 Case (iii): linear cointegration in a nonlinear VECM
{adjustwidth}
-2cm-2cm Refer to caption Refer to caption
@@
This case, which is depicted in the 3.4, entails that no trends are present in $\{y_{t}\}$, which is in fact a stationary process. The common trends are loaded entirely on $\{x_{t}\}$, and the cointegrating relationships between the elements of $x_{t}$ are unaffected by the sign of $y_{t}$, exactly as in literature on regime-switching nonlinear VECM models (see the text following (LABEL:nlvecm) above). For this reason, a model configured as in case (iii) is appropriate only when the univariate behaviour of $\{y_{t}\}$ appears stationary, in which case there will be frequent switches of ‘regime’, particularly if the mean of $y_{t}$ is near zero.
@@ 4 Representation theory and asymptotics
To facilitate the exposition, we shall initially suppose the data to be generated by a canonical CKSVAR, i.e. DGP holds, and that there are no deterministic trends in cases (i) and (ii), i.e. DET holds. Theorems 4.14.3 below are stated under these assumptions: the minor modifications required when DGP is replaced by DGP are given in the subsequent remarks, while a general treatment of deterministic trends follows in 4.4. To avoid the need to specify how some of the quantities below should be defined when $k=1$, we shall treat this as a special case of the model with $k=2$, in which $\phi_{2}^{+}=\phi_{2}^{-}=0$ and $\Phi_{2}^{x}=0$; and thus henceforth $k\geq 2$, unless otherwise stated. Because of the overlap between the arguments used to analyse the CKSVAR in each case, it will be occasionally necessary to redefine objects that have already appeared in the discussion of another case. While we have endeavoured to keep such notational conflicts to a minimum, and explicitly indicated wherever these arise, the reader is advised to treat each of the following three subsections, and the accompanying subsections of B where the proofs of the theorems appear, as independent of each other.
@@ 4.1 Case (i): regulated cointegration
Recalling 3.1, we have the familiar factorisation
@@
where $\alpha^{+},\beta^{+}\in\mathbb{R}^{p\times r}$ have rank $r$. As we show below, $\beta^{+}$ spans the cointegrating space on $\mathscr{Z}^{+}=\mathbb{R}_{+}\times\mathbb{R}^{p-1}$ (recall 3.2), so that $(y_{t}^{+},x_{t})\sim I^{\ast}(1)$ but $\beta^{+\mathsf{T}}(y_{t}^{+},x_{t})\sim I^{\ast}(0)$. Moreover, since $y_{t}^{-}\sim I^{\ast}(0)$, it follows that $\beta^{+\mathsf{T}}(y_{t},x_{t})\sim I^{\ast}(0)$, so that the columns of $\beta^{+}$ are globally cointegrating vectors.
@@
denote its joint spectral radius (JSR; e.g. Jungers, 2009, Defn. 1.1), for $\mathcal{A}^{t}\coloneqq\{\prod_{s=1}^{t}M_{s}\mid M_{s}\in\mathcal{A}\}$ the set of $t$-fold products of matrices in ${\cal A}$ [^10] For a further discussion of the JSR, and references to the literature on methods for numerically approximating it, see Duffy et al. (2023, Sec. 4.1). Let $z_{t}\coloneqq(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$.
@@
[…]
3
@@
    for $U_{0}(\lambda)\coloneqq\Gamma^{+}(1){\cal Z}_{0}+U(\lambda)$,
@@
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via 2.1. (See D.)
(ii). The contrast with a linear cointegrated VAR is marked. $Y$ is now a regulated Brownian motion, which also enters into other components of $X$. Indeed, as noted in 3.3 above, some components of $X$ may themselves be regulated BMs.
(iii). Part (i) of the theorem is proved by obtaining a nonlinear VAR representation for $\xi_{t}^{+}$, $\Delta z_{t}$ and $y_{t}^{-}$, whose companion form can be expressed in terms of the matrices $\{\boldsymbol{F}_{\delta}\mid\delta\in[0,1]\}$ (B.1). Since the parameters of that VAR depend on $y_{t}^{+}$, these processes cannot be stationary, but CO(i).2 ensures that the system is sufficiently ‘constrained’ that they will be $I^{\ast}(0)$. A necessary but not sufficient condition for CO(i).2 is that $\boldsymbol{F}_{0}$ and $\boldsymbol{F}_{1}$ have all their eigenvalues inside the unit circle. (This is implied by CVAR: see E.)
@@
(v). In deriving the weak limit of $n^{-1/2}y_{\lfloor n\lambda\rfloor}$, a key step is to obtain a univariate representation for $y_{t}^{+}$ as a regulated process. The proof of 4.1 shows that it is possible to write $y_{t}^{+}-\kappa_{1}y_{t}^{-}=w_{t}$ for a certain series $\{w_{t}\}$: the role of CO(i).3 is to ensure that this equation is solved uniquely by taking $y_{t}^{+}=[w_{t}]_{+}$ and $y_{t}^{-}=\kappa_{1}^{-1}[w_{t}]_{-}$. It is possible that this condition may be redundant: as shown in E, it is implied by our other assumptions if either $k=1$ or $p=1$.
@@
In the univariate ($p=1$) model (LABEL:univariate-case), case (i) with $r=0$ corresponds to a model in which $c=0$, $\sum_{i=1}^{k}\phi_{i}^{+}=1$ and $\phi^{-}(\lambda)$ has all its roots outside the unit circle, so that
@@
may be loosely regarded as an autoregressive model with a unit root regime (since $\pi^{+}=0$) and a stationary regime (though the model technically has $2^{k}$ distinct autoregressive regimes). 4.1 implies that $y_{t}^{-}\sim I^{\ast}(0)$, and
@@
on $D[0,1]$, where $K(\lambda)=\gamma^{+}(1)^{-1}U(\lambda)$, for $\gamma^{+}(1)=1-\sum_{i=1}^{k-1}\gamma_{i}^{+}$. $Y$ is thus a regulated Brownian motion. (LABEL:univariate-regulated) extends the results of Liu et al. (2011, Thm. 3.1) and Gao et al. (2013) from a first- to a higher-order autoregressive setting. It also agrees with the limit theory developed in Bykhovskaya and Duffy (2024, Thm. 3.2), when their censored dynamic Tobit model (in which $\phi_{i}^{-}=0$ for all $i\in\{1,\ldots,k\}$) is specialised to one with an exact unit root and no intercept.

### v4

@@ 3.3.2 Case (ii): kinked cointegration
Refer to caption Refer to caption
Refer to caption Refer to caption
@@
This case is perhaps more reminiscent of linear cointegration, which it accommodates as a special case. Here $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$ trace out ‘half’ subspaces of the same dimension $q$, but they need not be parallel, giving rise to a kink in $\mathscr{M}$ at the origin, with $\mathscr{M}$ itself being a linear cone. In general, $(Y,X)$ will follow a kinked Brownian motion driven by $p$ linear BMs, whose loadings, and the associated $r$ cointegrating relations, vary with the sign of $Y$. In the top panel of Figure 3.3, $y_{t}$ and $x_{t}$ are cointegrated, but with distinct cointegrating vectors $\beta^{+}=(1,-1)^{\mathsf{T}}$ or $\beta^{-}=(\tfrac{1}{2},-1)^{\mathsf{T}}$ applying on $\mathscr{Z}^{+}$ or $\mathscr{Z}^{-}$, i.e. when $y_{t}$ is positive or negative. We have a kind of ‘threshold cointegration’, with the movement of $y_{t}$ across zero causing the model to switch between distinct cointegrating spaces (cf. Saikkonen and Choi, 2004). This switch is reflected in the trajectories of $y_{t}$ and $x_{t}$: when $y_{t}\geq 0$, the two series move together approximately one-for-one; whereas if $y_{t}\leq 0$, $y_{t}$ changes by two units for every one-unit change in $x_{t}$, in the long run. Relative to case (i), the trajectories of $\{y_{t}\}$ will now much more closely resemble those of a linear unit root process, and $\{y_{t}\}$ will accordingly tend to spend long stretches in both the positive and negative regions, with no tendency to revert to one or the other.
The bottom panel of Figure 3.3 presents an important special case where the cointegrating relationships are such that $\{x_{t}\}$ behaves like an $I^{\ast}(0)$ process when $y_{t}<0$, but cointegrates with $y_{t}$ when $y_{t}>0$; the limiting process $X$ will thus be a censored Brownian motion.
@@
This model (with $\overline{\pi}_{t}=\overline{\pi}=0$) may be rendered as
@@
$r_{t}^{\ast}$ $=r_{t-1}^{\ast}+\eta_{t}.$ (2.6)
Then similarly to (3.6), we have from the corresponding CKSVAR for $(i_{t},\pi_{t})$, as given in (2.7) above, that
@@
where $\gamma^{-1}\tau_{\mu}=\theta(1-\mu)(1-\theta\gamma)^{-1}\leq 0$, with strict inequality unless $\mu=1$. Thus $\operatorname{rk}\Pi^{+}=\operatorname{rk}\Pi^{-}=1$, and (by Theorem 4.2 below), there is a common stochastic trend present in both regimes – but when $i_{t}>0$ this loads only on $i_{t}$, and the cointegrating vector is $\beta^{+}=(0,1)^{\mathsf{T}}$. If unconventional policy is as effective as conventional policy ($\mu=1$), this holds also when $i_{t}<0$; otherwise the trend is shared by both $i_{t}$ and $\pi_{t}$, and their cointegrating vector in the negative regime is $\beta^{-}=(\gamma^{-1}\tau_{\mu},1)^{\mathsf{T}}$. Qualitatively, the behaviour of the series is as plotted for $(y_{t},x_{t})$ in the second row of Figure 3.3 (with $y_{t}=-i_{t}$ and $x_{t}=-\pi_{t}$, i.e. with the signs reversed); in the limit, $n^{-1/2}\pi_{\lfloor n\lambda\rfloor}$ will be a censored BM (from above).
To account for this in economic terms, recall that (2.4a) and (2.5a) imply that when $i_{t}>0$, the central bank is able to fully stabilise inflation, in the sense that $\pi_{t}=\overline{\pi}_{t}+(1-\gamma\theta)^{-1}\varepsilon_{t}=(1-\gamma\theta)^{-1}\varepsilon_{t}$, since $\overline{\pi}_{t}$ is constant and normalised to zero. Thus $\pi_{t}\mathbf{1}\{i_{t}>0\}\sim I^{\ast}(0)$; whereas, $i_{t}^{+}\sim I^{\ast}(1)$, since it inherits the stochastic trend in the natural rate of interest. However, if the model solution requires $i_{t}<0$, then the ZLB constraint inhibits the operation of monetary policy (if $\mu<1$), and inflation begins to drift away from its target. That drift corresponds to the stochastic trend in $r_{t}^{\ast}$, which is thus present in both $i_{t}^{-}$ and $\pi_{t}$, and hence these series cointegrate.
By contrast, if monetary policy is not effectively constrained by the ZLB, then $\pi_{t}$ would remain $I^{\ast}(0)$, irrespective of the sign of $i_{t}$. Thus the long run behaviour of $\pi_{t}$ – whether it is $I^{\ast}(0)$, or whether it follows a stochastic trend (and so is $I^{\ast}(1)$) when interest rates are at the ZLB – here provides identifying information on the relative effectiveness of unconventional monetary policy, i.e. on whether the ZLB is ever a truly binding constraint on the central bank, just as it did in Example 2.1.
@@
Suppose that $\delta=0$ in (2.8), so the model is now described by (2.4b), (2.5b) and
@@
The inflation target thus remains time-varying, but there is no longer any feedback from past failures to hit that target. Then (3.6) simplifies to
@@
where $\varphi_{1}^{-1}\varphi_{\mu}=1+\frac{\theta(\gamma-1)(1-\mu)}{1-\theta\gamma}\in(0,1]$. Thus $i_{t}$ and $\pi_{t}$ are $I^{\ast}(1)$ and cointegrated, but with cointegrating relations $\beta^{+}=(1,-1)^{\mathsf{T}}$ and $\beta^{-}=(\varphi_{1}^{-1}\varphi_{\mu},-1)^{\mathsf{T}}$ that depend on the sign of $i_{t}$, unless $\mu=1$. Even though $i_{t}^{-}$ is unobserved, we can still distinguish between the cases $\mu=1$ and $\mu<1$ on the basis that, when $\mu<1$, the long-run variance of $\pi_{t}$ will differ depending on whether $i_{t}>0$ or $i_{t}=0$, as is evident in the behaviour of $x_{t}$ in the top panel of Figure 3.3.888Theorem 4.2 below implies that $n^{-1/2}\pi_{\lfloor n\lambda\rfloor}$ converges weakly to $\sigma_{\varepsilon}(\gamma-1)[1+(\varphi_{1}^{-1}\varphi_{\mu}-1)\mathbf{1}\{W(\lambda)<0\}]W(\lambda)$, for $\sigma_{\varepsilon}^{2}\coloneqq\mathbb{E}\varepsilon_{t}^{2}$ and $W$ a standard BM.
@@ 3.3.3 Case (iii): linear cointegration in a nonlinear VECM
Refer to caption Refer to caption
@@
This case, which is depicted in the Figure 3.4, entails that no trends are present in $\{y_{t}\}$, which is in fact a stationary process. The common trends are loaded entirely on $\{x_{t}\}$, and the cointegrating relationships between the elements of $x_{t}$ are unaffected by the sign of $y_{t}$, exactly as in literature on regime-switching nonlinear VECM models (see the text following (1.1) above). For this reason, a model configured as in case (iii) is appropriate only when the univariate behaviour of $\{y_{t}\}$ appears stationary, in which case there will be frequent switches of ‘regime’, particularly if the mean of $y_{t}$ is near zero.
@@ 4 Representation theory and asymptotics
To facilitate the exposition, we shall initially suppose the data to be generated by a canonical CKSVAR, i.e. DGP holds, and that there are no deterministic trends in cases (i) and (ii), i.e. DET holds. Theorems 4.14.3 below are stated under these assumptions: the minor modifications required when DGP is replaced by DGP are given in the subsequent remarks, while a general treatment of deterministic trends follows in Section 4.4. To avoid the need to specify how some of the quantities below should be defined when $k=1$, we shall treat this as a special case of the model with $k=2$, in which $\phi_{2}^{+}=\phi_{2}^{-}=0$ and $\Phi_{2}^{x}=0$; and thus henceforth $k\geq 2$, unless otherwise stated. Because of the overlap between the arguments used to analyse the CKSVAR in each case, it will be occasionally necessary to redefine objects that have already appeared in the discussion of another case. While we have endeavoured to keep such notational conflicts to a minimum, and explicitly indicated wherever these arise, the reader is advised to treat each of the following three subsections, and the accompanying subsections of Appendix B where the proofs of the theorems appear, as independent of each other.
Our proofs make repeated use of companion form representations of the VECM (3.2), formulated slightly differently for each of cases (i)–(iii). In this respect, the nearest analogue to our arguments, in the setting of a linear VAR, is provided by Hansen (2005).
@@ 4.1 Case (i): regulated cointegration
Recalling Table 3.1, we have the familiar factorisation
@@
where $\alpha^{+},\beta^{+}\in\mathbb{R}^{p\times r}$ have rank $r$. As we show below, $\beta^{+}$ spans the cointegrating space on $\mathscr{Z}^{+}=\mathbb{R}_{+}\times\mathbb{R}^{p-1}$ (recall Definition 3.2), so that $(y_{t}^{+},x_{t})\sim I^{\ast}(1)$ but $\beta^{+\mathsf{T}}(y_{t}^{+},x_{t})\sim I^{\ast}(0)$. Moreover, since $y_{t}^{-}\sim I^{\ast}(0)$, it follows that $\beta^{+\mathsf{T}}(y_{t},x_{t})\sim I^{\ast}(0)$, so that the columns of $\beta^{+}$ are globally cointegrating vectors.
@@
denote its joint spectral radius (JSR; e.g. Jungers, 2009, Defn. 1.1), for $\mathcal{A}^{t}\coloneqq\{\prod_{s=1}^{t}M_{s}\mid M_{s}\in\mathcal{A}\}$ the set of $t$-fold products of matrices in $\cal A$}. [^10] For a further discussion of the JSR, and references to the literature on methods for numerically approximating it, see Duffy et al. (2023, Sec. 4.1). Let $z_{t}\coloneqq(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$.
@@
      $n^{-1/2}z_{0}\overset{p}{\rightarrow}{\cal Z}_{0}=[\begin{smallmatrix}{\cal Y}_{0}\\ {\cal X}_{0}\end{smallmatrix}]\in\mathscr{M}$, where ${\cal Z}_{0}$ is non-random.
@@
    for $U_{0}(\lambda)\coloneqq\Gamma^{+}(1){\cal Z}_{0}+U(\lambda)$, jointly with (3.4),
@@
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via Proposition 2.1. (This obtains because the derived canonical variables then themselves follow a canonical CKSVAR that satisfies the conditions of the theorem; see Appendix D.)
(ii). The contrast with a linear cointegrated VAR is marked. $Y$ is now a regulated Brownian motion, which also enters into other components of $X$. Indeed, as noted in Section 3.3 above, some components of $X$ may themselves be regulated BMs.
(iii). Part (i) of the theorem is proved by obtaining a nonlinear VAR representation for $\xi_{t}^{+}$, $\Delta z_{t}$ and $y_{t}^{-}$, whose companion form can be expressed in terms of the matrices $\{\boldsymbol{F}_{\delta}\mid\delta\in[0,1]\}$ (Lemma B.1). Since the parameters of that VAR depend on $y_{t}^{+}$, these processes cannot be stationary, but CO(i).2 ensures that the system is sufficiently ‘constrained’ that they will be $I^{\ast}(0)$. A necessary but not sufficient condition for CO(i).2 is that $\boldsymbol{F}_{0}$ and $\boldsymbol{F}_{1}$ have all their eigenvalues inside the unit circle. (This is implied by CVAR: see Appendix E.)
@@
(v). In deriving the weak limit of $n^{-1/2}y_{\lfloor n\lambda\rfloor}$, a key step is to obtain a univariate representation for $y_{t}^{+}$ as a regulated process. The proof of Theorem 4.1 shows that it is possible to write $y_{t}^{+}-\kappa_{1}y_{t}^{-}=w_{t}$ for a certain series $\{w_{t}\}$: the role of CO(i).3 is to ensure that this equation is solved uniquely by taking $y_{t}^{+}=[w_{t}]_{+}$ and $y_{t}^{-}=\kappa_{1}^{-1}[w_{t}]_{-}$. It is possible that this condition may be redundant: as shown in Appendix E, it is implied by our other assumptions if either $k=1$ or $p=1$.
@@
In the univariate ($p=1$) model (2.10), case (i) with $r=0$ corresponds to a model in which $c=0$, $\sum_{i=1}^{k}\phi_{i}^{+}=1$ and $\phi^{-}(\lambda)$ has all its roots outside the unit circle, so that
@@
may be loosely regarded as an autoregressive model with a unit root regime (since $\pi^{+}=0$) and a stationary regime (though the model technically has $2^{k}$ distinct autoregressive regimes). Theorem 4.1 implies that $y_{t}^{-}\sim I^{\ast}(0)$, and
@@
on $D[0,1]$, where $K(\lambda)=\gamma^{+}(1)^{-1}U(\lambda)$, for $\gamma^{+}(1)=1-\sum_{i=1}^{k-1}\gamma_{i}^{+}$. $Y$ is thus a regulated Brownian motion. (4.5) extends the results of Liu et al. (2011, Thm. 3.1) and Gao et al. (2013) from a first- to a higher-order autoregressive setting. It also agrees with the limit theory developed in Bykhovskaya and Duffy (2024, Thm. 3.2), when their censored dynamic Tobit model (in which $\phi_{i}^{-}=0$ for all $i\in\{1,\ldots,k\}$) is specialised to one with an exact unit root and no intercept.